Prev Next

SW&IT Meeting Minutes 2018-12-03

Attendees: Andrew, Matthieu, Petr, Stephen, Yang

Outlook

Labor Reporting

Report auto-generated every, Do not edit this table, use entries in smartsheets.

20181105
LR20181105 - Bec, Matthieu 81.0
LR20181105 - Colson, Andrew 84.0
LR20181105 - Edmeades, David 32.0
LR20181105 - Sola, Igor 81.0
LR20181105 - Kubanek, Petr 65.5
LR20181105 - Hooper, Stephen 80.0
LR20181105 - Zhang, Yang 40.0
LR20181105 - Zhang, Xianyu 72.0

20181119
LR20181119 - Bec, Matthieu 64.0
LR20181119 - Colson, Andrew 78.0
LR20181119 - Edmeades, David 64.0
LR20181119 - Sola, Igor 53.0
LR20181119 - Kubanek, Petr 56.0
LR20181119 - Hooper, Stephen 64.0
LR20181119 - Zhang, Yang 40.0
LR20181119 - Zhang, Xianyu 0.0

Transition plan

Handover steps

  1. primary + secondary do self learning, gather information, read wiki
  2. primary organize teamviewer/zoom walk through with Kellee and secondary
  3. primary draft techtalk w. input from secondary
  4. techtalk review by Kellee
  5. primary does a team presentation

Please update this table to indicate progress.


Application(s) Primary Secondary Week
AzCam
AC (2)
Petr (2)
2
Weatherstation
AC (2)
Petr (2)
2
FACSUM,WEATHERSTATION
AC (2)
Matthieu (2)
4
ALH Igor(2)
Matthieu (2)
1
INDI Igor (2)
Matthieu (2)
3
Status Server Stephen (2)
Matthieu (2)
2
Telemetry cleanup Stephen (1) Petr 3
MODS
Yang (3)
Igor (1/2) 2
Telem-VIZ
Yang (1)
Stephen (1) 3
particle monitors AC Petr 4

News from management

  • PEPSI/pol high profile run in January: AGW3 & 4 must be ready
  • on-site interview Wed

Previous Action Items

  • Skill set list (DE)
  • Fill up team calendar with leave hours (DE)

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • network dropout - antenna monitoring?
  • utility building chiller ECS, Pepsi compressor - DanC
  • iLocater Single-Mode Data Fibers - CRA/CRB to Electronics Shop
  • BCU f/w upgrade documentation
  • revert simbad harvard-strasbourg /et/hosts hack from Friday
  • Dust monitor status
  • n505 mics x2
  • AO improvement: Luci kwds
  • AO cabinet switch upgrade - follow up
  • wifi summit upgrade (note: end-of-year deadline)
  • Tucson CR power overhaul plan (note: end-of-year deadline)
  • n507/obs upgrade (note: end-of-year deadline)
  • call out phone list - AC wants a cell

This week in github

Team reports

PK

  • patched PCS and GCS installed
  • It 7387: guiding with LBC works in NS, new problems emerged
  • patched GCS delivers same results as old GCS
  • C++11 telemetry and TCS ready for 2019A
  • optimized TCS code for C++11 (replaced auto_ptr with unique_ptr, removed not more needed Tr1 files)
next:

AC

  • Closed dome activities
    • closed loop on dx using cal unit
    • created new reconstructor on DX for testing
  • Investigate dx-lalas safety officer gui log messages (maybe from sx not being powered on)
  • RA/DEC decimals in argos_lan_satellite_avoidance_gui
  • Aarb State label flickering color in argos_laser_safety_officer_gui
  • Dichroic state not correctly populated in argos_dx_arbitrator_engineering_gui
  • AGW4 tested and ready to be shipped to mountain
  • Test ARGOS keywords being written to LUCI

YZ

  • Reading and working on MODS documents.
  • Attended the instruments meeting on Nov 29.
  • Working on issue #7285 from the instruments meeting (MODS1 data machine freeze problem).
  • Created Github repo for recent codes related to AO cravat project.
  • Will work on integrating some of Matthieu's JS9 hacks.
  • LBC guiding meeting on Dec 04.

ISS

Short week for me:

  • LUCI
    • Made small spread sheet on LUCI FITS AO/ARGOS keywords provenance. https://docs.google.com/spreadsheets/d/1jV-q7h8YTrFYzd_B9IKgAmJfxMi0NrcLnHc5URwEDt0/edit#gid=0
    • Save AFC subwindows and display them in a python program on demand. The program will be a standalone program that listens to subwindow file events, when it receives one such event it will rename and save the file locally and alert the python program to upload the new image. This software will help sciops to debug the luci AFC backend.
  • OT
    • Look at possible use of proper motion to display target in TPE. Put on hold till Michelle finds out more about it.
Next week
Finish the LUCI work.

MB

  • OC, WPC, management
  • OVMS+TT monthly
  • SHARK-NIR monthly
  • S/W onsite interview, candidate #2 phone interview
  • discuss Piston-Reconstruction Experiment (P-REx) w/ Perera Saavidra
  • AO headers investigation
  • svn2git mods
  • f29 play + download
next:
  • automate weekly report heading section
  • revamp software wiki main page
  • science archive updates from IA2
  • delta-tau training
  • iif/indi for new instruments
  • s/w recruting, new scheme

SH

  • continued working on Wifi summary
  • added information to http://info
  • AO cabinet switches tested, failed, reverted.
  • continuing mountain network woes.
next:
  • iLocater network and cabinets
  • AO cabinet switch configuration (try to feed 100Mbps in the lab).
  • Microwave information addition to info
  • Computers/cards for obs replacement
  • Finish Microwave stuff
  • Save history for info.
Topic revision: r17 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback